OpenStudy (ny,ny):

The parking lot has spaces for 450 cars. The ratio of spaces for full sized cars to compact cars is 11 to 4. Write a system of equations to solve the problem. How many spaces are for full sized cars? how many spaces are for compact cars?

OpenStudy (anonymous):

F + C = 450 F/C = 11/4

OpenStudy (ny,ny):

ok, yeah I had those equations down too. f + c = 450 f=450-c --------------- f/c = 11/4 450 - c/ c = 11/4 450 = 2c x 11/4 am i going the right way?

OpenStudy (ny,ny):

i feel like im doing something wrong...

OpenStudy (anonymous):

You were doing ok up until that last equation. It might be because you did not put 450 - c in parentheses because both those terms are over c, so 450 - c = (11c)/4 Then, 450 = (15c)/4 1800 = 15c c = 120 so f=330

OpenStudy (ny,ny):

how did 11c become 15c?

OpenStudy (anonymous):

I added c to each side in the form of 4c/4 because the right side had a 4 in the denominator.

OpenStudy (ny,ny):

ok. could you explain that out a little better please? I'm kind of confused.

OpenStudy (ny,ny):

sorry about that.

OpenStudy (anonymous):

\[450 - c = \frac{ 11c }{ 4 }\]so\[450 = \frac{ 11c }{ 4 } + c\]so\[450 = \frac{ 11c }{ 4 } + \frac{ 4c }{ 4 }\]

OpenStudy (ny,ny):

oh ok. that helps a lot better. thank you.

OpenStudy (anonymous):

you're welcome!